what is the surface area of a room that has a length of 89 inches, width 90 inches, and height 96 inches? please show the work for this problem
Answer by cleomenius(959)

That would be the formula for the surface area for a Right Rectangular Prism.
S = 2(ab + ac + bc)
a = 89
b = 90
c = 96
ab = 8010
ac = 8544
bc = 8640
8010 + 8544 + 8640 = 25194
SA = 2 (25194) = 50388 sq inches
Cleomenius.